OpenStudy (anonymous):

Please help! Select the pair of equations whose graphs are perpendicular. A. y = 2x – 7 x + 2y = 3 B. x + 6y = 8 y = 2x – 8 C. 2y = –3x + 5 2x + 3y = 4 D. 5x – 8y = 9 12x – 5y = 7

OpenStudy (anonymous):

find the slopes of each pair. if the slopes of the line multiply to give -1 then those lines are perpendicular

OpenStudy (anonymous):

The answer?

OpenStudy (anonymous):

another way is to put the lines into standard form (ax + by = c). then multiply the x coefficients and add that to the product of the y coefficients. If you get 0 then the lines are pependicular.

OpenStudy (anonymous):

so which one

OpenStudy (anonymous):

\[\begin{matrix}y=2x-7 \\ x+2y=3\end{matrix} \Rightarrow \begin{matrix}-2x+y=-7 \\ x+2y=3\end{matrix} \Rightarrow (-2)(1)+(1)(2)=0 \Rightarrow \perp\]

OpenStudy (anonymous):

a,b,c,or d

OpenStudy (anonymous):

a right?

OpenStudy (anonymous):

Are you here for answers, to learn or both?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

just tring to make sure if its right cause I just learned how to solve it so

OpenStudy (anonymous):

So how did solve?